OpenStudy (anonymous):

what is the solution of the system? Y=-8x-2 and Y=-6x+4 (-3,22) (-8/3,1/22) (-14/3,6/22) (3/14,-6/14) will give medal for correct answer

pooja195 (pooja195):

Use susbstituion :) plug equation 1 into equation 2

pooja195 (pooja195):

\[\huge -6x+4=-8x-2\]

OpenStudy (welshfella):

you can either solve the system or plug in the given values to see if they fit eg for first one y = 22 -8x - 2 = -8(-3) - 2 = 24 - 2 = 22 they both come to same value so thay fit for first equation Now try the same values for y = -6x + 4 and see if they fit

OpenStudy (anonymous):

i think its a

OpenStudy (welshfella):

r u sure ?

OpenStudy (welshfella):

be confident

OpenStudy (anonymous):

im not sure

OpenStudy (welshfella):

what does -6(-3) + 4 work out to?

OpenStudy (anonymous):

22

OpenStudy (welshfella):

so?

OpenStudy (anonymous):

it is a

OpenStudy (welshfella):

right

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

could you help me with this one how many solutions does the system have ? 12y=-3x+20 y=-1/4x+5/3

OpenStudy (welshfella):

yw pooja195 's method was quicker than mine though as both -6x+4 and -8x - 2 are equal to y they are equal to each other so you only needed to plug x = 3 into his equation to see if LHS = RHS

OpenStudy (welshfella):

the way to do these is to multiply one equation by a number so that the terms in y or x have the same coefficient ( the coefficient is the number in front of the x or y)

OpenStudy (welshfella):

so as we have 12y in equation 1 and y in equation 2 we multiply the second one by 12 equation 1 12y = - 3x + 20 equation 2 12y = -3x + 20 - the same equation so how many solution s does this have?

OpenStudy (anonymous):

1

OpenStudy (welshfella):

No - Hint; you can put any value of x into the equation and work out a value of y - in fact many many values of x...

OpenStudy (welshfella):

so theres not just 1 value theres ?

OpenStudy (anonymous):

infinte

OpenStudy (welshfella):

right infinite solutions

OpenStudy (anonymous):

thank you so much